OpenStudy (anonymous):

Solve the equation for y: 2x + 3(y + 9) = 17 ? CAN SOMEONE HELP ME ?

OpenStudy (anonymous):

first distribute 3 to the y and 9

OpenStudy (anonymous):

first distribute. 2x+3y+27=17 then subtract the 27 from 17 2x+3y=-17 then move the 2x over to the right 3y=-2x-17 then divide the 3 which makes it y=-2x-17/3

OpenStudy (anonymous):

2x+3y+27=17

OpenStudy (anonymous):

SO 2X +3Y + 27

OpenStudy (anonymous):

if you're solving for y, you wanna get it alone..

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

so then you get, 3y=-2x+27

OpenStudy (anonymous):

yeah that kid dwade is right

OpenStudy (anonymous):

IM LOST

OpenStudy (anonymous):

@Dwade03 he gave it to you

OpenStudy (anonymous):

NO THATS NOT NONE OF MY CHOICES